OpenStudy (inkyvoyd):

The length of a cube is increased by a factor of 3. By what factor does the surface area of the cube increase by?

OpenStudy (inkyvoyd):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

dunno the answer dude haha

OpenStudy (inkyvoyd):

:( does anyone know how to do this?

OpenStudy (bahrom7893):

v = x^3 V = (3x)^3

OpenStudy (bahrom7893):

how much bigger is V?

OpenStudy (inkyvoyd):

V=27x^3?

OpenStudy (bahrom7893):

oh wait you asked for surface area, not volume lol sorry.

OpenStudy (inkyvoyd):

so 27 then?

OpenStudy (bahrom7893):

Same concept though. A cube has 6 sides. a = 6(x^2)

OpenStudy (bahrom7893):

what am i thinking, it's 9

OpenStudy (inkyvoyd):

so 6 then?

OpenStudy (inkyvoyd):

why 9? a cube has 6 sides

OpenStudy (bahrom7893):

instead of squaring x, square 3x.

OpenStudy (inkyvoyd):

ok, so 3x^2=3x^2? so 3x^2 then?

OpenStudy (bahrom7893):

Okay, since a cube has 6 sides, and the area of each side is x^2: a = x^2 + x^2 + x^2 + x^2 + x^2 + x^2

OpenStudy (inkyvoyd):

what is x

OpenStudy (bahrom7893):

x is the side

OpenStudy (inkyvoyd):

oh okay

OpenStudy (inkyvoyd):

so a=6x^2 but surface area of a square is x^2 so it's 6 isn't it

OpenStudy (xapproachesinfinity):

haha, this is entertaining somehow!

OpenStudy (inkyvoyd):

@xapproachesinfinity after the fourth time you lose faith in humanity

OpenStudy (inkyvoyd):

anyways I ran out of time so I just left it blank. Apparently the answer was 9.

OpenStudy (inkyvoyd):

thanks for helping though @bahrom7893

OpenStudy (xapproachesinfinity):

well bahrom was right?!

OpenStudy (inkyvoyd):

shouldn't it be 6 though?

OpenStudy (xapproachesinfinity):

yes if you want it to be! if you make your rule haha

OpenStudy (inkyvoyd):

?? math is never my rule

OpenStudy (bibby):

Why do you lose hope for humanity just because people couldn't/didn't answer your vague math problem

OpenStudy (xapproachesinfinity):

i said that if you make your rule that will be correct! if you are following someone else's rule you have to abide haha

OpenStudy (inkyvoyd):

I guess @xapproachesinfinity ... too bad the rules are already made

OpenStudy (inkyvoyd):

ps:the answer is 3 actually

OpenStudy (xapproachesinfinity):

well that's true

OpenStudy (adrynicoleb):

Kitten